Determine the density in g/L of a 5.770L sample of nitrous oxide at a pressure of 0.7450atm and a temperature of 355.0K. Show your work.
Answer #1

According to PV = nRT

PV = (wt/mol.wt)*RT

P = (wt/volume)(1/mol.wt)*RT

P = d*(1/mol.wt)*RT

d = (P*mol.wt)/RT

P =0.745 atm. , T = 355 K, mol.wt = 44, R = 0.0821 lit*atm*mol^-1*K^-1

d = (0.745*44)/(0.0821*355)

d = 1.125 gm/lt

d = (